OpenStudy (anonymous):

use the intercept to graph the equation x+3y=6

OpenStudy (anonymous):

There are two types of intercepts: x-intercepts ( when y=0 ) and y-intercepts ( when x=0 ) To use intercepts for the equation x + 3y = 6: X-intercepts: Since you know that y will be 0 with x intercepts, PLUG IN 0 FOR Y: x + 3y = 6 x + 3(0) = 6 x = 6 Therefore, the x intercept is 6. The same process is used for finding y intercepts, except that this time x is always 0 ( not y ). Once you find both the x and y intercepts, plot the two points and draw the line. :D